# Bridging Binarization: Causal Inference with Dichotomized Continuous Exposures

### v4

@@ 3.2 Visualizing the MTPs
Figure 1: Densities of exposures AA, $\tilde{A}_{1}$, and $\tilde{A}_{0}$ by different values of $W$. The graph on the left shows the densities when $W=0$ and graph on the right shows the densities when $W=1$. The green line represents the cut off value of $A=6$. The gray line is the density of the observed $A$. The red dotted line is the density of $\tilde{A}_{1}$. The blue dotted line is the density of $\tilde{A}_{0}$.
@@ 3.3 Results
Table 1: Simulation results for estimating $\Psi_{BATE}$. The true value of the estimand is 201.806.
@@
Table 2: Simulation results for estimating $\Psi_{CAB}$. The true value of the estimand 90.872.
Using the methods described in Section 2.6, we calculated estimates for the target parameters and the standard errors of the estimates. What bears repeating is that the only information the estimators require about the exposure $A$ is the value of the binarized exposure $T$. One may assume that, in order to estimate the expected outcomes under the MTPs, one must first estimate the full densities of $\tilde{A}_{1}$ and $\tilde{A}_{0}$. However, through the equivalence with the binarized ATE, the relevant quantity one must estimate is instead the conditional probability $\pi_{\mathcal{A}}(w)=Pr(T=1|W=w)=Pr(A\geq 6|W=w)$ for $w\in\{0,1\}$, which is easier to estimate.
@@
4 Application: Oil/Gas Well Exposure and Birth Outcomes in California, 2007-2015
@@
To demonstrate how to think about binarization in an applied setting, we revisit the questions analyzed by Tran et al. investigating the effect of residential proximity of oil/gas wells during pregnancy on birth outcomes in the state of California from 2006—2015 [32]. The authors found a significant association between high production volume from oil and gas wells and small gestational age (SGA) in rural and urban areas. In response to“a growing body of research shows direct health impacts from proximity to oil extraction,” California Senate passed a law creating Health Protection Zones from oil and gas wells [33]. The law creates Health Protection Zones 3,200 feet ($\approx 1$ kilometer) around homes, schools, hospitals, and other sensitive locations starting January 1, 2023. To evaluate the effect of such a law on the probability of SGA in newborns in California based on the 2007-2015 data, we considered a binarized exposure, where infants born to households within 1 km of oil and gas wells are considered “untreated,”, and infants born to households outside of 1km of an active oil/gas well are considered “treated.”
@@
Our population of interest is babies born in California between 2007 and 2015 where the residential address of the pregnant parent was within 10km of an active oil/gas well during the period of gestation, as in Tran et al - we followed the same exclusion criteria as in Figure 1. We consider the same infant and maternal confounding factors as in Tran et al. — specifically, we adjust for include infant sex, month (categorical) of birth, year of birth, maternal age in years, maternal race/ethnicity, maternal educational attainment, Kotelchuk index of prenatal care, and parity. All study protocols were approved by the Institutional Review Board of the CA Department of Public Health (#13-05-z).
@@
Let $\mathbf{W}$ be the vector of covariates, $D$ be the distance in meters to the closest well, and $Y$ be the indicator of low birth weight (defined as a birthweight of <2500g<2500g [32]). Some parents may live within 10km of multiple wells - for the purposes of this analysis, we chose the distance to the closest well as our exposure of interest. We then binarize this exposure at 1km (assigning individuals with a distance $D\leq 1$km to $T=0$ and a distance D>1D>1km $T=1$. Our estimands of interest are $\Psi_{BATE}=E[E[Y|T=1,\mathbf{W}]]-E[E[Y|T=0,\mathbf{W}]]$ and $\Psi_{CAB}=E[E[Y|T=1,\mathbf{W}]]-E[E[Y|\mathbf{W}]]$.
@@
To estimate these quantities, we use the AIPW estimator presented in A.2. We estimate the outcome regression $E[Y|T,\mathbf{W}]$ using a logistic regression estimator: $\widehat{m}(T,\mathbf{W})=\text{expit}(\widehat{\beta_{0}}+T\widehat{\beta_{T}}+\mathbf{W}\widehat{\mathbf{\beta}}_{\mathbf{W}}+T\mathbf{W}\widehat{\mathbf{\beta}}_{interact}$). To estimate the propensity score ($P(D>1km)$), we use a logistic regression estimator: $\widehat{\pi_{T}}(w)=\text{expit}\left(\widehat{\beta}_{0}+\mathbf{W}\widehat{\beta}_{\mathbf{W}}\right)$. If either of these estimators are consistent for the true function, then the estimates will be consistent. To estimate the standard deviation, we implement the nonparametric bootstrap. The results are presented in Table 3.
@@
Table 3: Estimates for $\Psi_{BATE}$ and $\Psi_{CAB}$ when binarizing distance to closest active oil/gas well at 1km.
The key identifying assumption underlying our binarization is that, conditional on observed covariates, the distribution of distances to the nearest well among those who would be “treated” under the law (i.e., living more than 1 km from an active well) is proportional to the distribution observed in the current world. In our notation,
@@
Figure 2: Histograms for the observed distance to the nearest active well to households with pregnant Hispanic people, aged 25–29, less than high school educated with a Kotelchuk index of adequate and nulliparous, who gave birth in September 2010 to male babies. The histogram on the left is the observed distribution. The histogram on the right is the observed histogram after imposing a cut-off at 1 km — this is approximately what we assume the distribution of distance to nearest active well would look like after implementing the Health Protection Zone.
It may help to consider what this means in terms of the California law. When the law is enacted, households do not relocate. Instead, the wells themselves are removed (or relocated beyond the buffer). Thus, the intervention acts by eliminating wells within 1 km of residences and perhaps relocating them outside of 1 km. Under the binarization assumption, the distances to the nearest well beyond 1 km remain unchanged. In reality, however, removing wells near households will often increase the distances to the nearest active well for other households, since the nearest active well may now be farther away. Therefore, the assumption is not literally true, though the degree of violation depends on the density and spatial configuration of wells. When wells are sparse, the shift could be substantial; when wells are dense, the effect may be negligible.
Finally, it is important to emphasize that the binarization assumption is not the only requirement for causal validity. A causal interpretation of this analysis still depends on the the assumption in Section 2.4 holding for our analysis. For example, we must have no unmeasured unconfounding. Despite controlling for maternal demographics, education, prenatal care, and birth timing, unmeasured socioeconomic or environmental factors (some of which were included in Tran et al.) could bias results. In addition, our parametric modeling choices introduce assumptions. Logistic regressions for outcome regressions and propensity scores are unlikely to be correctly specified in this context, which can induce bias. More flexible machine learning methods with cross-validation could mitigate these concerns, but pursuing that is beyond the scope of this paper. We therefore interpret our results as illustrative of how binarization can be applied, rather than definitive evidence of health effects.
@@
It may seem surprising that the BATE estimate here is positive, implying that a policy moving all wells outside of 1 km of households would increase the probability of low birth weight compared to a policy moving all wells within 1 km of households. There are two reasons our binarized estimand may be understating the protective effect of the policy. The first is that, as explained above, if the health harms of proximity decay smoothly with distance, then removing nearby wells should reduce exposure more than what is captured by a strict 1 km cut-off. Put differently, the real law is expected to do at least as well as, and likely better than, the scenario we analyze. On the other hand, if effects beyond 1 km are truly negligible, then the difference is immaterial, and the binarization assumption becomes a reasonable approximation for causal inference. The second is the exposure we consider, distance to closest well, may not fully capture the health harms of allowing active wells to be built within 1 km of households. For example, some individuals may live within 1 km of several wells, and some active wells may have higher levels of production than others. For example, the authors of the Tran et al. paper found that households exposed to high production volume had an average of 32 active wells within 1 km of the household. Thus, considering only the closest well may not capture the full health harms pregnant people are exposed to prior to the buffer zone being implemented. The fact that the estimates are very close to 0 lend further credence to this idea, suggesting that the distance to closest well may not be a good proxy for the amount of harm an individual may experience living within 1km of active wells.
@@ 5 Discussion
We also introduce a new estimand, the causal attributable effect of binarization, which contrasts the expected outcome under one of the MTPs associated with binarization with the expected outcome in the observed world. This estimand requires weaker identification assumptions that are implied by the assumptions required to identify the binarized ATE – thus, a scenario in which researchers could estimate the BATE is also a scenario in which they could estimate the CAB. We argue that this estimand answers a more relevant causal question as it compares the counterfactual world under one MTP (the “treated” counterfactual) to the observed world (which is akin to a “control” counterfactual). As this estimand cannot be calculated using standard methods developed for the ATE, we also provide details for how to implement three different estimators of the causal attributable effect of binarization as well as how to estimate their standard errors.
@@
Although the binarized ATE and the causal attributable effect of binarization may be easy to estimate, they may not be relevant causal quantities depending on the specific research context. Ultimately, whether binarization should be employed as detailed in this paper depends on the expected behavior of individuals after implementing the intervention of interest. For example, suppose researchers were interested in estimating the effect of a policy that limits the amount of energy households can use by fining households that consume more energy than a certain threshold. In this case, the assumption of preservation of relative self-selection is not very plausible, as one could reason that households above the threshold would change their energy usage to just barely meet the new limits, as there is no incentive to reduce energy consumption any further. Conversely, households that already use less energy than the threshold will likely make no change to their energy use. Therefore, the resulting distribution of household energy usage would have a “piling up” of households at the imposed threshold with the distribution below the threshold looking the same as that in the observed world. Therefore, in this case, it would likely not make sense to binarize household energy usage based on the proposed cut-off.
@@ A.1 Regression Estimator
$\hat{Y}=\hat{\beta_{0}}+T\hat{\beta}_{T}+\widecheck{W}^{\prime}\hat{\beta}_{W}+T\widecheck{W}^{\prime}\hat{\beta}_{interact}$
@@
$\widehat{\Psi}_{BATE}^{reg}$ $=\hat{\beta}_{T}$ (4)
[…]
@@
$\widehat{\mathbb{V}}\left(\widehat{\Psi}_{BATE}^{reg}\right)$ $=\begin{bmatrix}0&0&0&1&0&0\\ \end{bmatrix}\mathbb{V}(\hat{\beta})\begin{bmatrix}0\\ 0\\ 0\\ 1\\ 0\\ 0\\ \end{bmatrix}$ (7)
$\widehat{\mathbb{V}}\left(\widehat{\Psi}_{CAB}^{reg}\right)$ $=\begin{cases}\begin{bmatrix}-\hat{\beta_{T}}&-\hat{\beta}_{interact}&0&(1-\overline{T})&0&-\overline{T\widecheck{W}^{\prime}}\end{bmatrix}\widehat{\mathbb{V}}(\hat{\beta})\begin{bmatrix}-\hat{\beta_{T}}\\ -\hat{\beta}_{interact}\\ 0\\ (1-\overline{T})\\ 0\\ -\overline{T\widecheck{W}}\end{bmatrix}\text{ for }t=1\\ \begin{bmatrix}-\hat{\beta_{T}}&-\hat{\beta}_{interact}&0&-\overline{T}&0&-\overline{T\widecheck{W}^{\prime}}\end{bmatrix}\widehat{\mathbb{V}}(\hat{\beta})\begin{bmatrix}-\hat{\beta_{T}}\\ -\hat{\beta}_{interact}\\ 0\\ -\overline{T}\\ 0\\ -\overline{T\widecheck{W}}\end{bmatrix}\text{ for }t=0\end{cases}$ (8)

### v5

@@ 3.2 Visualizing the MTPs
Figure 1: Densities of exposures AA, $\tilde{A}_{1}$, and $\tilde{A}_{0}$ by different values of $W$. The graph on the left shows the densities when $W=0$ and graph on the right shows the densities when $W=1$. The green line represents the cut-off value of $A=6$. The gray line is the density of the observed $A$. The red dotted line is the density of $\tilde{A}_{1}$. The blue dotted line is the density of $\tilde{A}_{0}$.
@@ 3.3 Simulation Results
Table 1: Simulation results for estimating $\Psi_{BATE}$. The true value of the estimand is 201.806. The “Estimate” and “Bias” columns contains the average of the BATE estimates and bias over the 5000 replicates, respectively. “Est SD” column contains the average of the estimated standard deviations over the 5000 replicates. Finally, the “Sim SE” column contains the standard error of the estimates over the 5000 replicates.
@@
Table 2: Simulation results for estimating $\Psi_{CAB}$. The true value of the estimand is 90.872. The “Estimate” and “Bias” columns contains the average of the CAB estimates and bias over the 5000 replicates, respectively. “Est SD” column contains the average of the estimated standard deviations over the 5000 replicates. Finally, the “Sim SE” column contains the standard error of the estimates over the 5000 replicates.
Using the methods described in Appendix A, we calculated estimates for the target parameters and the standard errors of the estimates. What bears repeating is that the only information the estimators require about the exposure $A$ is the value of the binarized exposure $T$. One may assume that, in order to estimate the expected outcomes under the MTPs, one must first estimate the full densities of $\tilde{A}_{1}$ and $\tilde{A}_{0}$. However, through the equivalence with the binarized ATE, the relevant quantity one must estimate is instead the conditional probability $\pi_{\mathcal{A}}(w)=Pr(T=1|W=w)=Pr(A\geq 6|W=w)$ for $w\in\{0,1\}$, which is easier to estimate.
@@
4 Application: Oil/Gas Well Exposure and Birth Outcomes in California, 2007–2015
@@
To demonstrate how to think about binarization in an applied setting, we revisit the questions analyzed by Tran et al. investigating the effect of residential proximity of oil/gas wells during pregnancy on birth outcomes in the state of California from 2006–2015 [31]. The authors found a significant association between high production volume from oil and gas wells and small gestational age (SGA) in rural and urban areas. In response to“a growing body of research shows direct health impacts from proximity to oil extraction,” California Senate passed a law creating Health Protection Zones from oil and gas wells starting January 1, 2023 [2]. Health Protection Zones are 3,200 feet ($\approx 1$ kilometer) buffers surrounding homes, schools, hospitals, and other sensitive locations. To evaluate the effect of such a law on the probability of SGA in newborns in California based on the 2007-2015 data, we considered a binarized exposure, where infants born to households within 1km of oil and gas wells are considered “untreated,” and infants born to households outside of 1km of an active oil/gas well are considered “treated.”
@@
Our population of interest is infants born in California between 2007 and 2015 where the residential address of the pregnant parent was within 10km of an active oil/gas well during the period of gestation, as in Tran et al - we followed the same exclusion criteria as in Figure 1. We consider the same infant and maternal confounding factors as in Tran et al.; specifically, we adjust for infant sex, month (categorical) of birth, year of birth, maternal age in years, maternal race/ethnicity, maternal educational attainment, Kotelchuk index of prenatal care, and parity. All study protocols were approved by the Institutional Review Board of the CA Department of Public Health (#13-05-z).
@@
Let $\mathbf{W}$ be the vector of covariates, $D$ be the distance in meters to the closest well, and $Y$ be the indicator of low birth weight (defined as a birthweight of <2500g<2500g [31]). Some parents may live within 10km of multiple wells — for the purposes of this analysis, we chose the distance to the closest well as our exposure of interest. We then binarize this exposure at 1km (assigning individuals with a distance $D\leq 1$km to $T=0$ and a distance D>1D>1km $T=1$. Our estimands of interest are $\Psi_{BATE}=E[E[Y|T=1,\mathbf{W}]]-E[E[Y|T=0,\mathbf{W}]]$ and $\Psi_{CAB}=E[E[Y|T=1,\mathbf{W}]]-E[E[Y|\mathbf{W}]]$.
@@
To estimate these quantities, we use the AIPW estimator presented in Appendix A.2. We estimate the outcome regression $E[Y|T,\mathbf{W}]$ using a logistic regression estimator: $\widehat{m}(T,\mathbf{W})=\text{expit}(\widehat{\beta_{0}}+T\widehat{\beta_{T}}+\mathbf{W}\widehat{\mathbf{\beta}}_{\mathbf{W}}+T\mathbf{W}\widehat{\mathbf{\beta}}_{interact}$). To estimate the propensity score ($Pr(D>1\text{km}|\mathbf{W})$), we use a logistic regression estimator: $\widehat{\pi_{T}}(w)=\text{expit}\left(\widehat{\beta}_{0}+\mathbf{W}\widehat{\beta}_{\mathbf{W}}\right)$. If either of these estimators are consistent for the true function, then the estimates will be consistent. To estimate the standard deviation, we implement the nonparametric bootstrap. The results are presented in Table 3.
@@
Table 3: Estimates for $\Psi_{BATE}$ and $\Psi_{CAB}$ and the corresponding standard deviations when binarizing distance to closest active oil/gas well at 1km.
The key identifying assumption underlying our analysis is that, conditional on observed covariates, the distribution of distances to the nearest well among those who would be “treated” under the law (i.e., living more than 1 km from an active well) is proportional to the distribution observed in the current world. In our notation,
@@
Figure 2: Histograms for the observed distance to the nearest active well to households with pregnant Hispanic people, aged 25–29, less than high school educated with a Kotelchuk index of adequate and nulliparous, who gave birth in September 2010 to male babies. The histogram on the left is the observed distribution. The histogram on the right is the observed histogram after imposing a cut-off at 1 km; this is approximately what we assume the distribution of distance to nearest active well would look like after implementing the Health Protection Zones.
It may help to consider what this means in terms of the California law. When the law is enacted, households do not relocate — instead, the wells themselves are removed (or relocated beyond the buffer). Thus, the intervention acts by eliminating wells within 1 km of residences and perhaps relocating them outside of 1 km. Under the binarization assumption, the distances to the nearest well beyond 1 km remain unchanged. In reality, however, removing wells near households will often increase the distances to the nearest active well for other households, since the nearest active well may now be farther away. Therefore, the assumption is not literally true, though the degree of violation depends on the density and spatial configuration of wells. When wells are sparse, the shift could be substantial; when wells are dense, the effect may be negligible.
Finally, it is important to emphasize that the binarization assumption is not the only requirement for causal validity. A causal interpretation of this analysis still depends on the assumptions in Section 2.4 holding for our analysis. For example, we must have no unmeasured confounding. Despite controlling for maternal demographics, education, prenatal care, and birth timing, unmeasured socioeconomic or environmental factors (some of which were included in Tran et al.) could bias results. In addition, our parametric modeling choices introduce assumptions. Logistic regressions for outcome regressions and propensity scores are unlikely to be correctly specified in this context, which can induce bias. More flexible machine learning methods with cross-validation could mitigate these concerns, but pursuing that is beyond the scope of this paper. We therefore interpret our results as illustrative of how binarization can be applied, rather than definitive evidence of health effects.
@@
It may seem surprising that the BATE estimate here is positive, implying that a policy moving all wells outside of 1 km of households would increase the probability of low birth weight compared to a policy moving all wells within 1 km of households. There are two reasons our binarized estimand may be understating the protective effect of the policy. The first is that, as explained above, if the health harms of proximity decay smoothly with distance, then removing nearby wells should reduce exposure more than what is captured by a strict 1 km cut-off. Put differently, the real law is expected to do at least as well as, and likely better than, the scenario we analyze. On the other hand, if effects beyond 1 km are truly negligible, then the difference is immaterial, and the binarization assumption becomes a reasonable approximation for causal inference. The second is the exposure we consider, distance to closest well, may not fully capture the health harms of allowing active wells to be built within 1 km of households. For example, some individuals may live within 1 km of several wells, and some active wells may have higher levels of production than others. For example, the authors of the Tran et al. paper found that households exposed to high production volume had an average of 32 active wells within 1 km of the household. Thus, considering only the closest well may not capture the full health harms pregnant people are exposed to prior to the buffer zone being implemented. The fact that the estimates are very close to 0 lends further credence to this idea, suggesting that the distance to closest well may not be a good proxy for the amount of harm an individual may experience living within 1km of active wells.
@@ 5 Discussion
We also introduce a new estimand, the causal attributable effect of binarization (CAB), which contrasts the expected outcome under one of the MTPs associated with binarization with the expected outcome in the observed world. Identification of this estimand requires weaker assumptions, and these assumptions are implied by the assumptions required to identify the BATE — thus, a scenario in which researchers could estimate the BATE is also a scenario in which they could estimate the CAB. We argue that this estimand answers a more relevant causal question as it compares the counterfactual world under one MTP (the “treated” counterfactual) to the observed world (which is akin to a “control” counterfactual). As this estimand cannot be calculated using standard methods developed for the ATE, we also provide details for how to implement four different estimators of the causal attributable effect of binarization as well as how to estimate their standard errors.
@@
Although the BATE and CAB may be easier to estimate, they may not be relevant causal quantities depending on the specific research context. Ultimately, whether binarization should be employed as detailed in this paper depends on the expected behavior of individuals after implementing the intervention of interest. For example, suppose researchers were interested in estimating the effect of a policy that limits the amount of energy households use by fining households that consume more energy than a certain threshold. In this case, the assumption of preservation of relative self-selection is not very plausible, as one could reason that households above the threshold would change their energy usage to just barely meet the new limits, as there is no incentive to reduce energy consumption any further. Conversely, households that already use less energy than the threshold will likely make no change to their energy use. Therefore, the resulting distribution of household energy usage would have a “piling up” of households at the imposed threshold with the distribution below the threshold looking the same as that in the observed world. Therefore, in this case, it would likely not make sense to binarize household energy usage based on the proposed cut-off.
@@ Acknowledgments
Funding information
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. The research was partially supported by funds from the NIEHS Superfund Research Program P42ES004705.
Author contributions
All authors have accepted responsibility for the entire content of this manuscript and approved its submission.
Conflict of interest
Authors state no conflict of interest.
Ethical approval
The research related to human use has been complied with all the relevant national regulations, institutional policies and in accordance the tenets of the Helsinki Declaration, and has been approved by the authors’ institutional review board or equivalent committee.
Data availability statement
The code generating the simulation study data and analyzing the simulation and applied data are available at https://github.com/kaitlynjlee/bridging_binarization. The births data that support the findings of this study are available from the California Department of Public Health, but restrictions apply to the availability of these data, which are governed under an IRB protocol to protect the confidentiality of participants. The well data are a subset of the California Geologic Energy Management Division’s (CalGEM) All Wells dataset, available at https://www.conservation.ca.gov/calgem/maps/Pages/GISMapping2.aspx.
@@ A.1 Regression Estimator
$\hat{Y}=\hat{\beta_{0}}+T\hat{\beta}_{T}+\widecheck{W}^{\prime}\hat{\beta}_{W}+T\widecheck{W}^{\prime}\hat{\beta}_{interact},$
@@
$\widehat{\Psi}_{BATE}^{reg}$ $=\hat{\beta}_{T}$
$\widehat{\Psi}_{CAB}^{reg}$ $=\begin{cases}(1-\overline{T})\hat{\beta}_{T}-\overline{T\widecheck{W}^{\prime}}\hat{\beta}_{interact}\text{ for }t=1\\ (-\overline{T})\hat{\beta}_{T}-\overline{T\widecheck{W}^{\prime}}\hat{\beta}_{interact}\text{ for }t=0\\ \end{cases},$
@@
$\widehat{\mathbb{V}}\left(\widehat{\Psi}_{BATE}^{reg}\right)$ $=\begin{bmatrix}0&0&0&1&0&0\\ \end{bmatrix}\mathbb{V}(\hat{\beta})\begin{bmatrix}0\\ 0\\ 0\\ 1\\ 0\\ 0\\ \end{bmatrix}$
$\widehat{\mathbb{V}}\left(\widehat{\Psi}_{CAB}^{reg}\right)$ $=\begin{cases}\begin{bmatrix}-\hat{\beta_{T}}&-\hat{\beta}_{interact}&0&(1-\overline{T})&0&-\overline{T\widecheck{W}^{\prime}}\end{bmatrix}\widehat{\mathbb{V}}(\hat{\beta})\begin{bmatrix}-\hat{\beta_{T}}\\ -\hat{\beta}_{interact}\\ 0\\ (1-\overline{T})\\ 0\\ -\overline{T\widecheck{W}}\end{bmatrix}\text{ for }t=1\\ \begin{bmatrix}-\hat{\beta_{T}}&-\hat{\beta}_{interact}&0&-\overline{T}&0&-\overline{T\widecheck{W}^{\prime}}\end{bmatrix}\widehat{\mathbb{V}}(\hat{\beta})\begin{bmatrix}-\hat{\beta_{T}}\\ -\hat{\beta}_{interact}\\ 0\\ -\overline{T}\\ 0\\ -\overline{T\widecheck{W}}\end{bmatrix}\text{ for }t=0\end{cases},$